### v1

# LLM-based relevance assessment still can’t replace human relevance assessment

Abstract
The use of large language models (LLMs) for relevance assessment in information retrieval has gained significant attention, with recent studies suggesting that LLM-based judgments provide comparable evaluations to human judgments. Notably, based on TREC 2024 data, Upadhyay et al. [8] make a bold claim that LLM-based relevance assessments, such as those generated by the UMBRELA system, can fully replace traditional human relevance assessments in TREC-style evaluations. This paper critically examines this claim, highlighting practical and theoretical limitations that undermine the validity of this conclusion. First, we question whether the evidence provided by Upadhyay et al. [8] really supports their claim, particularly if a test collection is used as a benchmark for future improvements. Second, through a submission deliberately intended to do so, we demonstrate the ease with which automatic evaluation metrics can be subverted, showing that systems designed to exploit these evaluations can achieve artificially high scores. Theoretical challenges – such as the inherent narcissism of LLMs, the risk of overfitting to LLM-based metrics, and the potential degradation of future LLM performance – must be addressed before LLM-based relevance assessments can be considered a viable replacement for human judgments.
1 Introduction
Almost two years ago, Faggioli et al. [5] applied GPT-3.5 (text-davinci-003) to fully reassess runs submitted to the TREC 2021 Deep Learning track. They reported Kendall’s $\tau=0.86$ for human vs. LLM-based assessment on NDCG@10. A natural conclusion might be that LLMs could now replace humans for routine relevance assessment. Instead, Faggioli et al. [5] issue a warning. While recognizing the potential of LLMs to improve ranking and acknowledging their value as part of the relevance assessment process, they argue strongly against abandoning human assessment. They raise concerns about the potential for unknown biases that LLM-based assessments might introduce. They highlight the issue of circularity, where LLMs evaluate the outputs of other LLMs. Most importantly, their primary concern is that “LLMs are not people.” Since information retrieval systems are designed to serve human needs, their evaluation must ultimately reflect human judgment and preferences.
Recently, Upadhyay et al. [8] analyze the retrieval task results from the TREC 2024 RAG track. This retrieval task (or “R task”) mirrors a traditional TREC ad hoc retrieval task. Participating systems were tasked with executing 301 queries over the MS MARCO Segment V2.1 collection, producing a ranked set of 100 passages for each query (a “run”). They compare four procedures for assessing runs: 1) fully automatic assessments using the UMBRELA LLM-based relevance assessment tool [9]; 2) fully manual assessments using established TREC evaluation protocols for human judgments; 3) a hybrid method where UMBRELA filtered the set of passages to be judged; and 4) a hybrid method where humans refined UMBRELA’s assessments. In this paper, we focus on the first two procedures: fully automatic and fully manual assessments. Based on their analysis, Upadhyay et al. [8] conclude:
Our results suggest that automatically generated UMBRELA judgments can replace fully manual judgments to accurately capture run-level effectiveness. Surprisingly, we find that LLM assistance does not appear to increase correlation with fully manual assessments, suggesting that costs associated with human-in-the-loop processes do not bring obvious tangible benefits…. Our work validates the use of LLMs in academic TREC-style evaluations and provides the foundation for future studies.
We disagree. Not only does their reported correlation fail to provide stronger evidence than that of Faggioli et al. [5], but additional evidence from the track directly contradicts their conclusion. This evidence includes runs submitted by one of the authors’ research teams (WaterlooClarke), which were explicitly designed to subvert LLM-based relevance judgments by employing LLM-generated judgments as a final-stage ranker.
Faggioli et al. [5] already demonstrated a strong empirical correlation between manual judgments and LLM judgments, both in terms of inter-annotator agreement and leaderboard correlation, and many other work also observed this empirical correlation. However, after a detailed consideration of competing views, Faggioli et al. [5] concluded that there are too many theoretical concerns before human judgments can be replaced. These concerns, which remain critical to the discussion, have neither been addressed nor refuted in the work of Upadhyay et al. [8, 9].
Conceptually, there is no fundamental difference between an LLM-based relevance assessment and an LLM-based re-ranking method. Human relevance judgments are privileged precisely because they are created by humans, and only humans can provide a gold standard for the evaluation of usefulness. While employing LLMs to train and implement rankers can lead to substantial performance gains, these improvements risk being illusory if they fail to reflect human judgments. The observation that LLM-based relevance judgments closely mimic the outcomes of human relevance judgments suggests that these LLM assessments may themselves represent a strong ranking method, rather than a valid evaluation metric.
We acknowledge the value of work by Upadhyay et al. [9]. In particular, their use of a new collection with fresh queries, which guarantees that LLMs were not trained on this collection. Moreover, their work confirms that the correlation seen by Faggioli et al. [5] is not merely an artifact of training on the test collection. However, the warnings issued by Faggioli et al. [5] remain both valid and increasingly urgent, especially given the growing prevalence of LLM-based relevance assessments in information retrieval tasks.
[…]
Figure 1: Scatter plot extracted from Figure 3 of Upadhyay et al. [8] comparing manual assessment and automatic assessment. Each red dot represents the average performance of a run over all queries. The blue dots plot all run-query combination. The inset provides a closer view of the points in the green rectangle.
2 Differences among Top-Performing Systems
While Upadhyay et al. [8] report an overall Kendall’s $\tau=0.89$ for manual vs. automatic assessment, the correlation appears much weaker on the top-scoring runs, where agreement with manual judgments matters most. Although Kendall’s $\tau$ is a suitable metric for showing overall rank correlations on large-scale experiments, it is less informative for assessing progress at the top of the leaderboard. Demonstrating methodological advancements often involves reusing test collections with the goal of surpassing the best-performing systems. Therefore, identifying meaningful differences at the top of the leaderboard is critical for measuring significant progress. Based on the results reported by Upadhyay et al. [9], automatic judgments fail to demonstrate strong alignment with manual judgments for top-performing systems. This misalignment undermines the reliability of automatic assessments in distinguishing improvements among top performing systems.
Closer consideration of results from Upadhyay et al. [8] highlights specific discrepancies in top-performing system rankings. Figure 1 reproduces a scatter plot extracted from Figure 3 of that paper, showing the performance of submitted runs (red dots) under manual assessment (y-axis) vs. automatic assessment (x-axis).111 At the time of writing, Upadhyay et al. [8] do not provide a full public data release. They have generously provided limited access to their data for the purpose of confirming factual statements in this paper. Focusing on the top-performing systems in the inset of Figure 1, discrepancies between manual and automatic evaluations become evident. For example, the system ranked highest under automatic evaluation would only place fifth under manual evaluation. Conversely, the top system under manual evaluation ranks sixth under automatic evaluation. A particularly interesting is the case is the run circled in green. While it ranks fifth under automatic evaluation, it drops dramatically to 28th under manual evaluation.
These inconsistencies underscore a fundamental limitation of LLM-based assessments, such as those used by UMBRELA, in reliably identifying the best-performing systems. As a result, caution is warranted when using these methods for evaluating and validating progress in retrieval tasks. While these differences may be the attributed to statistical noise — only 27 queries were manually judged, and no error bars are provided — the evidence presented is insufficient to justify the replacement of human judgments. This concern is particularly important when a collection is intended for re-use. To substantiate claims of a novel system’s superiority over state-of-the-art methods, we must ensure that any observed improvements, measured through fully automatic judgments, are both meaningful and aligned with manual assessments. At present, such confidence is lacking.
3 Subverting Automatic Evaluation
When task relevance labels are generated entirely through a publicly known automatic process, such as UMBRELA, the evaluation metric becomes vulnerable to manipulation. For instance, a participant could pool the output of many rankers, apply the UMBRELA system to this pool, and submit the resulting relevance labels as a participating system for evaluation. This approach could theoretically achieve perfect scores across all metrics. Even if the specific LLM-based relevance assessment process includes undisclosed elements, such as the exact prompt or LLM used, participants could approximate the process to subvert automatic evaluations.
3.1 Empirical Demonstration of the Risk
The run circled in green in Figure 1 exemplifies this approach. Submitted by Clarke’s research group (WaterlooClarke) as run uwc1, this submission was deliberately designed to subvert the automatic evaluation process. They pooled the top 20 documents from 15 preliminary runs, spanning neural and traditional rankers, with and without query expansion. This pool was then judged by GPT-4o with the prompt described by Faggioli et al. [5]. Top-graded passages were subsequently judged pairwise following the procedure of Clarke et al. [3] but using LLM-based assessments, rather than crowdsourced human judgments.
The final ranking for uwc1 was determined using LLM-based preference judgments as the primary key, LLM-based relevance assessments as the secondary key, and the reciprocal rank fusion [4] of preliminary runs as the tertiary key. As intended, this deliberate attempt to manipulate the evaluation process led uwc1 to rank significantly higher under automatic assessment (5th) than under manual assessment (28th).
3.2 LLM-based Relevance Assessment as Re-Ranking
LLM-based relevance assessment is effectively just another form of LLM-based re-ranking. To employ an LLM-based relevance assessment tool as a LLM-based re-ranker, one starts with an initial ranking and we prompt an LLM to assign a score — expressed as a relevance grade — to each of its top-$k$ passages. These passages are then re-ranked according to their relevance grades, preserving their original order when grades are equal.
This concept is further illustrated by another run from WaterlooClarke, uwc2, which demonstrates that LLM-based relevance assessment operates as a pointwise re-ranking method. For uwc2, the group re-ranked the track’s baseline run using the prompt described by Arabzadeh and Clarke [2]. This re-ranking improves the baseline’s performance from 9th to 4th under manual assessments and from 7th to 3rd under automatic assessments.
4 Automatic Relevance Labels are not Gold Standards
The prompts used to elicit relevance grades from LLM-based assessment tools resemble instructions typically given to human assessors. However, this resemblance is superficial and we should not be fooled by it. Such prompts merely represent one of many possible ways an LLM-based re-ranking method might assign scores, akin to an a LLM-based point-wise ranker. Despite being commonly referred to as relevance assessments, these scores are not equivalent to the judgments produced by humans.
LLM-based relevance assessments cannot serve as a gold standard because they lack the grounding in human judgment necessary to evaluate the usefulness of retrieval systems. A true gold standard must originate from human assessments, as only humans can determine the relevance of information in a way that reflects real-world utility.
Faggioli et al. [5] raised concerns about the potential unknown biases inherent in LLM-based assessments. However, one clear and concerning bias is that LLM-based relevance assessments tend to favor LLM-based ranking systems. This bias has been observed in other contexts as well [6, 10, 7], where LLMs demonstrate a form of “narcissism,” disproportionately favoring outputs generated by similar models. Furthermore, Alaofi et al. [1] show that LLMs can be deceived through well-crafted prompt attacks embedded in content, leading them to incorrectly judge irrelevant text as relevant. These vulnerabilities highlight not only the susceptibility of LLM-based assessments to manipulation but also their inability to objectively evaluate diverse ranking approaches. Such biases and flaws further undermine the reliability of LLM-based assessments as a substitute for human judgments in critical tasks.
5 When Automatic Judgments become Useless
While the uwc1 run demonstrates how a bad actor can strategically subvert an evaluation experiment, it is reasonable to assume that most participants are well-intentioned. These participants are not merely competing to win but are contributing to the creation of reusable test collections that support the development of innovative systems. However, even without malicious intent, the next generation of information systems will likely incorporate the latest advancements in LLMs, including prompting for relevance. As a result, some ranking methods will inherently embed elements that mirror LLM-based relevance judgments. In a future evaluation experiment, it is plausible that even a well-intentioned participant could inadvertently undermine the evaluation process.
Looking ahead, we anticipate that future retrieval systems will increasingly rely on automatically generated training data to optimize machine learning components. Here, Goodhart’s law serves as a cautionary principle: “When a measure becomes a target, it ceases to be a good measure.” While the current observed correlation between manual and automatic assessment methods is strong, we predict that this correlation will degrade as developers incorporate LLM-based evaluation components into their systems. Over time, these systems risk becoming disconnected from the human judgments they are intended to serve. If the entire end-to-end experimental pipeline – from query formulation to relevance labeling – is fully automated, the evaluation process devolves into an LLM assessing its own assessments. The circularity feared by Faggioli et al. [5] is no longer a hypothetical concern; it has already begun to manifest in practice.
6 Conclusion
This paper raises serious concerns about the claims made by Upadhyay et al. [8], which presents a preliminary analysis of data from the retrieval (“R”) task of the TREC 2024 RAG Track. The author list of Upadhyay et al. [8] includes some of the most prominent experts in the area of information retrieval evaluation. Despite being preliminary, their conclusions strongly imply that LLM-based relevance assessment can replace human relevance assessment — a claim that does withstand scrutiny. Given the authority of the authors and the strength of their implied conclusions, there is a risk that these findings may gain widespread acceptance within the research community without sufficient critical consideration. Nearly two years ago, Faggioli et al. [5] reached the opposite conclusion based on similar evidence. Their concerns have still not been addressed.
Acknowledgments
We have discussed our concerns with some of the authors of Upadhyay et al. [8], and we appreciate their attention and feedback. They also generously provided us with access to their non-public data to confirm factual statements in this paper.